OpenStudy (anonymous):

What is the mass in grams for 1.52 moles of Li? I believe the answer is 10.5488 grams but would like to make sure that that is true.

OpenStudy (imtiaz7):

molar mass of li is 6.9410 so here we'll convert it to 1.52 moles of it so: 1.52 moles of li= 6.9410*1.52 so it equals to 1.05508 grams

OpenStudy (zpupster):

\[(\frac{ 1.52 moles}{ 1})(\frac{ 6.941 g }{ moles })=? g of Li\]

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

yup zpupster that is what i have

OpenStudy (anonymous):

amd Imitaz7 are you supposed to divide by 10 at the end?

OpenStudy (anonymous):

\[moles \text{ or moles?} \] \text{or moles} in latex